# Spectroscopic Characterization of Interstellar Object 3I/ATLAS: Water Ice in the Coma

### v1

@@ 4 Discussion
As the most abundant solid in protoplanetary disks, water ice plays a key role in planet formation (B. Bitsch & A. Johansen, 2016; J. Müller et al., 2021). In the Solar System, it likely contributed to the cores of giant planets (F. J. Ciesla, 2014) and is a major component of comets (M. F. A’Hearn, 2017; M. J. Mumma & S. B. Charnley, 2011). The detection of water ice in the coma of 3I/ATLAS suggests that it formed beyond the snowline of its parent system, and that water ice could be its major component, similar to comets in the solar system. The current brightening is consistent with activity driven by water-ice sublimation, but since no gas emission has yet been observed, there could also be activity driven by super-volatiles from the sub-surface such as CO or CO2. The H2O/CO ratio remains unknown, making future searches for CO and CO2 with ground- and space-based facilities essential.
@@
The physical structure of water ice records its formation conditions. Below 100 K, water forms thermodynamically unstable amorphous ice, which gradually converts to crystalline ice above 137 K (A. Kouchi et al., 1994). Due to limited SNR in the IRTF data, we cannot constrain the ice phase through spectral modeling. However, the water ice signature appears to be strengthening as the comet’s activity increases. IRTF spectra taken 10 days earlier set a 7% upper limit on coma ice (T. Kareta et al., 2025), while our observations suggest that water ice accounts for approximately 30% of the coma by area. During this period, 3I moved from =4.4 to =4.1 au and brightened significantly. The increased activity not only enhanced the comet’s brightness, improving the signal-to-noise ratio of the spectrum, but may also have promoted the ejection of additional icy grains. Additionally, imperfect telluric correction in the T. Kareta et al. (2025) spectrum may have limited their ability to accurately model the water-ice feature. Upcoming JWST/NIRSpec observations are expected to provide deeper insight into the ice properties and composition.
@@
Beyond the absorption feature near 2.0 m, the spectral slope of the continuum provides key insights into the dust composition of the coma. Several studies have reported optical spectral slopes for 3I/ATLAS using various instruments and observing conditions, with values ranging from 17% to 19% 4% per 1000 Å (C. Opitom et al., 2025; D. Z. Seligman et al., 2025). Similar large ranges of measured spectral slope is also reported among previous studies of 1I and 2I.
@@
Our measurement of = 10% per 1000 Å, derived from Gemini South/GMOS data taken on July 5, is slightly lower than this range. The observed variation in spectral slope could be due to using different solar analog stars. The near-infrared slope measured from IRTF/SpeX over the 0.9–1.5 m range is 3.060.21% per 1000Å, noticeably flatter than the optical slope, consistent with the presence of water ice in the coma.
@@
Despite being obtained 9 days apart with different instruments, the GMOS and SpeX datasets show excellent agreement in the overlapping 0.8–0.9 m region, suggesting temporal stability in the continuum shape and confirming the consistency of the optical–NIR transition. This comparison underscores the importance of coordinated, multi-wavelength observations in characterizing transient interstellar objects. Single-instrument datasets may be affected by instrumental or environmental biases, but combined optical and infrared spectra enable a more robust and comprehensive understanding of composition and activity.
@@ 5 Summary
The near-infrared spectrum shows a neutral to slightly bluish slope of % per 1000 Å over the 1.5-2.3 m range. The flattening beyond 1.5 m and the presence of a broad absorption band centered near 2.0 m are consistent with the presence of water ice grains 10 m in radius, contributing approximately 30% of the coma by area.
@@
The optical spectrum of 3I appears featureless and exhibits a moderately red slope consistent with D-type material of % per 1000 Å, (between 0.5-0.8 m) consistent with the slope range reported in earlier studies.
@@
The GMOS and SpeX datasets, taken with different instruments and separated by 9 days, display consistent spectral slopes in the overlapping 0.8–0.9 m range. This consistency indicates that no significant temporal evolution occurred in the continuum shape during that period.
Together, these results suggest that 3I/ATLAS hosts a coma containing large water ice grains, and that its dust continuum is stable over at least week-long timescales. The spectral characteristics further distinguish 3I from known ultrared trans-Neptunian objects and align it more closely with active Jupiter-family comets.

### v2

@@ 4 Discussion
As the most abundant solid in protoplanetary disks, water ice plays a key role in planet formation (B. Bitsch & A. Johansen, 2016; J. Müller et al., 2021). In the Solar System, it likely contributed to the cores of giant planets (F. J. Ciesla, 2014) and is a major component of comets (M. F. A’Hearn, 2017; M. J. Mumma & S. B. Charnley, 2011). The detection of water ice in the coma of 3I/ATLAS suggests that it formed beyond the snowline of its parent system, and that water ice could be its major component, similar to comets in the solar system. The derived abundance of 37% water ice by volume in the coma is consistent with the theoretical study predicting a water-ice mass fraction of 20–40% for 3I-like objects (M. J. Hopkins et al., 2025).
@@
The physical structure of water ice records its formation conditions. Below 100 K, water forms thermodynamically unstable amorphous ice, which gradually converts to crystalline ice above 137 K (A. Kouchi et al., 1994). Due to limited SNR in the IRTF data, we cannot constrain the ice phase through spectral modeling. IRTF spectra taken 10 days earlier set a 7% upper limit on coma ice (T. Kareta et al., 2025), while our observations suggest that water ice accounts for approximately 37% of the coma by volume. In addition, recent IRTF observations show that the water ice signature appears to be stronger as the comet’s activity increases (C. M. Lisse et al., 2025). The increased activity not only enhanced the comet’s brightness, improving the SNR of the spectrum, but may also have promoted the ejection of additional icy grains. The JWST/NIRSpec observations clearly show strong water-ice absorption features at 3.0 and 4.5 m (M. A. Cordiner et al., 2025). Detailed spectral modeling of these data can provide valuable information on the size, temperature, and impurities of the water-ice grains, and place strong constraints on the formation conditions of this ISO.
@@
Beyond the absorption feature near 2.0 m, the spectral slope of the continuum provides key insights into the dust composition of the coma. Several studies have reported optical spectral slopes for 3I/ATLAS using various instruments and observing conditions, with values ranging from 16% to 27% per 1000 Å (C. Opitom et al., 2025; D. Z. Seligman et al., 2025; T. H. Puzia et al., 2025; B. T. Bolin et al., 2025). Similar large ranges of measured spectral slope is also reported among previous studies of 1I and 2I.
@@
Our measurement of = 11% per 1000 Å, derived from Gemini South/GMOS data taken on July 5, is lower than this range. The observed variation in spectral slope could be due to using different solar analog stars. The near-infrared slope measured with IRTF/SpeX over the 0.8–2.3 m range is % per 1000 Å, significantly flatter than the optical slope, and consistent with the presence of water ice in the coma.
@@
Despite being obtained 9 days apart with different instruments, the GMOS and SpeX datasets are statistically consistent in the overlapping 0.7–0.9 m region, indicating temporal stability in the continuum shape and confirming the reliability of the optical-NIR transition. This comparison underscores the importance of coordinated, multi-wavelength observations in characterizing transient interstellar objects. Single-instrument datasets may be affected by instrumental or environmental biases, but combined optical and infrared spectra enable a more robust and comprehensive understanding of composition and activity.
@@ 5 Summary
The optical spectrum of 3I/ATLAS, obtained with GMOS, is featureless and exhibits a moderately red slope of % per 1000 Å between 0.5 and 0.8 m. This spectral behavior is comparable to D-type asteroids and active comets in the Solar System, but markedly different from the ultrared objects in the trans-Neptunian region. The measurement is consistent with slopes reported in earlier studies and supports the interpretation that 3I’s dust is similar in reflectance properties to Solar System cometary material rather than the ultrared TNO population.
@@
The near-infrared spectrum from IRTF/SpeX displays a neutral to slightly bluish slope between 1.7 and 2.3 m, together with a broad 11% absorption band centered near 2.0 m. These features are well reproduced by a spectral model consisting of amorphous carbon mixed with 1.3 m-sized water ice grains, with water ice contributing about 37% of the coma volume. The flatter continuum relative to the optical, combined with the detection of the 2.0 m absorption, provides strong evidence that abundant water ice is present in the coma.
@@
The GMOS and SpeX datasets, taken with different instruments and separated by nine days, display consistent spectral slopes in the overlapping 0.7–0.9 m region. This agreement indicates that no significant short-term spectral evolution occurred during that period, the overall scattering properties of the coma remained stable over timescales of at least one week.
Our detection of H2O ice in 3I/ATLAS at 4 au with IRTF, together with the strong CO2 outgassing revealed by JWST and SPHEREx, supports the hypothesis that this object formed in a cold, volatile-rich region of its parent protoplanetary disk. These observations also demonstrate that such volatile compositions can be preserved through ejection and maintained during the subsequent interstellar journey.